HETAV PATEL

Grade 11,

If A1,A2,A3,....is in AP such that, A1+A5+A10+A15+A20+A24=225, then, A1+A2+A3+...A23+A24 is equal to?

If A1,A2,A3,....is in AP such that, A1+A5+A10+A15+A20+A24=225, then, A1+A2+A3+...A23+A24 is equal to?

Grade:11

2 Answers

Vikas TU
7 years ago
A1,A2,A3,....is in AP
the write in common terms of a and d.
.i.e.
A1 = a
A2  =a+d
A3 =a+2d
.
.
A24 = a+23d
summing all we get,  4(6a+69d) = > 4*225 = 900 perhaps.
krunal
7 years ago
u can also solve in such way,
(A1+A24) +(A5+A20) +(A10+A15)=225.
THEREFORE 3(A1+A24)=225
(A1+A24)=75.
NOW SUM=24/12(A1+A24)
=900